Consider the production function Q = 2(KL)0.5

a)     What is the marginal product of labour and capital

b)     What is the marginal rate of technical substitution of labor for capital

 c)     What is the elasticity of substitution at a point K = 1, L = 1 if we increase K by one unit?


 

1
Expert's answer
2021-05-16T17:50:04-0400

(a) "Q = 2(KL)^{0.5}"

"Q = 2 K^{0.5} L^{0.5}"

Marginal product of labor = Differentiate the above production function with respect to L

"\frac{dQ}{dL}= 0.5 × 2 K^{0.5 }L^{0.5-1}"

 "\frac{dQ}{dL}= 0.5 × 2 K^{0.5 }L^{-0.5}"

"\frac{dQ}{dL}=\frac {K^{0.5 }}{L^{0.5}}"

Marginal product of labor = K0.5 / L0.5


"Q = 2 K^{0.5 }L^{0.5}"

Marginal product of capital = Differentiate the above production function with respect to K

"\frac{dQ}{dK} = 0.5 × 2 × K^{0.5-1} L^{0.5}"

"\frac{dQ}{dK }= 0.5 ×2 × K^{-0.5 }L^{0.5}"

"dQ/dK =\frac{ L^{0.5}} { K^{0.5}}"

Marginal product of capital = L0.5 / K0.5

 

(b)marginal rate of technical substitution of labor for capital =change in capital/chang in labor

="\frac{∆K}{∆L}"

"\frac{L^{0.5} K^{0.5}}{K^{0.5} L^{0.5}}"

"=\frac{K}{L}"

(c) Elasticity of substitution σ is given by

"σ=\frac{d ln(k/l)}{d ln(|MRTS|)}"


"=\frac{d ln(k/l)}{d ln(|MRTS|)}=1"

"MRTS=\frac{k}{l}"

Taking log on both sides

"ln\frac{k}{l}=ln({|MRTS|})"

Taking derivative both sides we get

"d ln\frac{k}{l}=d ln(|MRTS|)"

"σ=\frac{d ln(\frac{k}{l})}{d ln(|MRTS|)}=1"